OpenStudy (anonymous):

A man flies a small airplane from Fargo to Bismarck, North Dakota --- a distance of 180 miles. Because he is flying into a head wind, the trip takes him 2 hours. On the way back, the wind is still blowing at the same speed, so the return trip takes only 1 hour 12 minutes. What is his speed in still air, and how fast is the wind blowing?

OpenStudy (shinalcantara):

let Vp=velocity of the plane Vw=velocity of the wind Vp-Vw=180miles/2hrs ----eqn 1 Vp=Vw+90mph ----eqn 2 Vp+Vw=180miles/1.2hrs Vp+Vw=150mph ------eqn 3 substitute eqn 2 to 3 Vw+90mph+Vw=150mph 2Vw=150-90 Vw=30mph substitute Vw to eqn 3 Vp+30mph=150mph Vp=150-30 Vp=120

OpenStudy (anonymous):

thank you!!!

OpenStudy (shinalcantara):

yw